### v2

@@ Abstract
We introduce SPAFormer, an innovative model designed to overcome the combinatorial explosion challenge in the 3D Part Assembly (3D-PA) task. This task requires accurate prediction of each part’s pose and shape in sequential steps, and as the number of parts increases, the possible assembly combinations increase exponentially, leading to a combinatorial explosion that severely hinders the efficacy of 3D-PA. SPAFormer addresses this problem by leveraging weak constraints from assembly sequences, effectively reducing the solution space’s complexity. Since assembly part sequences convey construction rules similar to sentences being structured through words, our model explores both parallel and autoregressive generation.It further enhances assembly through knowledge enhancement strategies that utilize the attributes of parts and their sequence information, enabling it to capture the inherent assembly pattern and relationships among sequentially ordered parts. We also construct a more challenging benchmark named PartNet-Assembly covering 21 varied categories to more comprehensively validate the effectiveness of SPAFormer. Extensive experiments demonstrate the superior generalization capabilities of SPAFormer, particularly with multi-tasking and in scenarios requiring long-horizon assembly. Codes and model weights will be released at https://github.com/xuboshen/SPAFormer.
Keywords:
Object assembly 3D point cloud Transformers
@@ 1 Introduction
Figure 1: Illustration of the combinatorial explosion challenge inherent in the assembly process. Specifically: (a) the number of constituent parts increases when the target object for assembly becomes more complex. (b) For an object composed of $n$ parts, where we assume each part can occupy one of $m$ discrete positions, the potential combinations of these parts grow at an extraordinary rate, exceeding $O(m^{n})$ in complexity.
@@
There has been a strong anticipation for agents to autonomously assemble parts into an intricate structural object [12, 13, 18, 45, 4]. To achieve this goal, 3D part assembly (3D-PA) is essential [13], which requires the ability of an agent to identify, reorient and connect geometric parts in the 3D space, ultimately forming a desired object. Such capability empowers an agent to undertake a variety of applications in real-world scenarios, ranging from assembling IKEA furniture [45] to crafting designs in Computer-Aided Design (CAD) environments [4].
@@
Due to its critical importance, the 3D-PA task has attracted increasing attention from researchers. Traditionally, many approaches [13, 3, 52] achieve target object assembly by relying solely on part-level information or basic structural knowledge (e.g., object category label) without additional constrains. We consider these approaches as condition-free approaches. However, the 3D-PA task encounters a substantial obstacle known as “combinatorial explosion” (depicted in Figure 1), where the number of possible combinations rises beyond exponentially with increasing parts. This phenomenon poses a significant challenge and makes condition-free approaches less effective in achieving accurate part assembly, particularly as the assembly sequence lengthens.
@@
To address the combinatorial explosion challenge, several studies [33, 26, 25, 3, 53, 24] have leveraged strong visual priors of the object, such as point clouds [26] or RGB images [25, 44], to inform the assembly process (illustrated in Figure 2 (b)). These visual-conditioned methods enable the foresight of an object’s final form and overall appearance to guide the assembly process. However, the acquisition of such detailed visual priors incurs significant costs, reducing the practicality of these approaches in real-world applications. As an alternative, recent initiatives explore a more balanced way by utilizing structural prior knowledge. This includes the adoption of a hierarchical structure graph [40, 31, 11, 16] or employing peg-hole joints on parts [24] for guidance (illustrated in Figure 2 (c)). While this strategy decreases the need for detailed visual priors and is generally less burdensome, the reliance on structural priors still implies an additional expense compared to condition-free approaches. Given the limitations of previous works, we propose that leveraging the assembly sequences represents a more viable solution (illustrated in Figure 2 (d)), which is advantageous due to its weak constrains: assembly sequences are readily obtainable from diverse sources, including instructional manuals [45, 44], the analysis of physical disassembly process [41, 28, 38], or through the prediction of assembly sequences themselves [29]. Similar to how the sequence of words in a sentence dictates the meaning, assembly sequences convey the rules for part construction. Consequently, recognizing patterns within these sequences can effectively benefit 3D-PA, offering an efficient method to address combinatorial explosion.
@@
In this work, we propose Sequential 3D Part Assembly with Transformers (SPAFormer), a novel transformer-based model to address the combinatorial explosion challenge inherent in the 3D-PA task. SPAFormer conditioned on the assembly sequence enhances the 3D geometric understanding through reasoning relations among object parts. Specifically, we explore the effectiveness of various potential transformer paradigms for 3D-PA, including parallel generator and autoregressive generator. Moreover, we employ knowledge enhancement strategies to enrich the model with part attributes and assembly sequence insights, by incorporating critical features such as part symmetry (e.g., the four legs of a table), assembly order, and the relative positioning of parts. To comprehensively evaluate our model, we expand the scope of the traditional benchmark [13], enriching it from the limited 3 categories (chair, table, lamp) to a wider variety of 21 object categories that range from structured furniture (e.g., chairs, tables) to everyday items (e.g., earphones, clocks). Our findings reveal that SPAFormer not only significantly outperforms existing methods that are condition-free or sequence-conditioned, it also achieves competitive performance comparable to approaches employing visual priors.
@@
    Innovative Framework: We propose SPAFormer, a transformer-based model for object assembly with sequential 3D parts. Our model particularly leverages sequential part information, and incorporates knowledge enhancement strategies to significantly improve the assembly performance.
@@
    Generalization of SPAFormer: SPAFormer shows superior generalization in object assembly from three crucial perspectives: a) category-specific, enabling it to handle various objects in the same category; b) multi-task, showcasing its versatility across diverse object categories using the same model; and c) long-horizon, proving its ability in managing complex assembly tasks with numerous parts.
@@
    A More Comprehensive Benchmark: To facilitate a thorough evaluation of different models, we introduce an extensive benchmark named PartNet-Assembly, covering up to 21 object categories and providing a broad spectrum of object assembly tasks for the first time.
Figure 2: Illustration of comparisons on different assembly constraints. The term "condition-free" refers to assembling without any predefined constraints, where the circle Refer to caption represent a part. In contrast to strong and costly conditions such as "visual priors" or "structure priors", in this paper, we propose to adopt "assembly sequence" which offers a reasonable solution to facilitate the assembly process.
@@ 2 Related Works
Assembly-based 3D Modeling. The assembly-based 3D modeling involves creating structural 3D shapes by integrating information from individual parts. Prior works [1, 17, 15, 9] have explored to construct diverse 3D shapes from a given set of part candidates with precise language semantics. With the development of generative methods, some works [48, 10, 31, 22, 11] focus on leveraging generative models, such as variational autoencoders [49, 21], to generate 3D shapes. Although generative models produce aethetic 3D shapes, they are not able to deal with assembling true objects in real-world. Many works focus on fracture object assemblies [27, 2, 35, 47, 20, 14], where the model aggregates broken pieces into an object. However, the application of restoring broken objects is limited in furniture assemblies or CAD modeling. Different from previous tasks, the autonomous 3D object assembly [13] is more beneficial for robotic assembly [19, 18, 12, 51, 4] where the parts maintains semantics like IKEA assembly.
@@
Autonomous 3D Object Assembly. In the context of 3D object assembly, the objective is to accurately predict the translation and rotation parameters of individual parts to reconstruct the object’s shape. Many studies [13, 52, 46, 3] try to tackle this task by simply utilizing the part information and category-level priors. Notably, Huang et al. [13] employ graph neural networks to facilitate the understanding of interrelations between parts. Additionally, expanding data through the inclusion of fractured objects, Zhang et al. [53] demonstrate significant improvements in assembling chairs. Despite these advancements, the problem of combinatorial explosion persists. To mitigate this issue, various prior knowledge on object has been provided as extra information, including ground-truth object image [25], peg-hole joints on parts [24] and object point clouds [26]. However, these constraints may be costly and less feasible in real-world applications. Instead, advancements in sequence-conditioned part assembly, such as RGL [33], have demonstrated the potential of using weakly-constrained assembly sequence for exploring free-form assembly. Building upon these insights, our work introduces a transformer-based framework designed to leverage the advantages offered by assembly sequences in the 3D-PA task.
@@
Positional Encoding in Transformers. Positional encoding facilitates transformers to interpret position information, as evidenced across diverse range of research domains [50, 5, 30, 23, 42, 6, 8]. Traditionally, the literature consists of two primary encodings: absolute and relative. Absolute positional encoding [43] enables transformers to differentiate words through position-dependent embeddings. Nonetheless, it exhibits limitations in representing the order of words relative to one another [36]. To address this issue, relative positional encoding aims to better characterize the relationships between adjacent words. A notable implementation is the RoPE [37], which rotates the input embedding to effectively extract relative positioning. In the 3D-PA task, Li et al. [25] and Narayan et al. [33] attempt to differentiate identical parts using positional encodings, but the unexpected similar features lead to their suboptimal results. Considering these works, we propose variants of positional encodings to identify assembly patterns among sequential parts.
@@ 3 Methodology
Given a set of $N$ object parts $\mathcal{P}=\{p_{i}\}_{i=1}^{N}$, the part assembly task aims to predict their poses $\{y_{i}\}_{i=1}^{N}\subset{\rm SE}(3)$, where $p_{i}$ denotes the 3D point clouds of the $i$-th part and ${\rm SE(3)}$ includes all translation and rotation transformations. The resulting assembled 3D shape $S=\bigcup_{i}y_{i}(p_{i})$ is formulated by transforming each part-level point cloud $p_{i}$ using its corresponding pose $y_{i}$. During the assembly process, the assembly chain $\mathcal{C}=\{(p_{i},p_{j})|p_{i}\preceq p_{j},\forall i\leq j\}$ is imposed on the set $\mathcal{P}$, where $p_{i}\preceq p_{j}$ denotes that part $p_{i}$ is assembled before $p_{j}$. The goal of our work is sequential part assembly, which predicts the poses $\{y_{i}\}_{i=1}^{N}$ of $\mathcal{P}$ guided by the predefined assembly chain $\mathcal{C}$.
The overall framework of our proposed SPAFormer is illustrated in Figure 3. In the following sections, we first introduce our knowledge enhancement strategies to effectively incorporate part attributes and relation information in Section 3.1. Then we discuss generator variants for better modeling part relationship in Section 3.2. These strategies includes order encoding (OEnc) and relation encoding (REnc) which aim to capture assembly sequence patterns, and symmetry encoding (SEnc) which focus on enriching part features with symmetry information. In Section 3.3, we introduce our training process.
Figure 3: Illustration of overall end-to-end framework of SPAFormer. (a) The shared 3D backbone extracts geometry feature of individual parts, follows by (b) knowledge enhancement of part features, which incorporates symmetry, order and relation information into part features through positional encodings, then generates poses by either (c1) parallel generator, which generates poses of all parts at once, or (c2) autoregressive generator, which decodes poses of parts according to assembly sequences step by step.
@@ 3.1 Knowledge Enhancement for Parts
Due to the data deficiency, which can lead to inferior results when using the data-hungry transformers, we employ several strategies to enhance the model’s understanding of parts. These strategies include order encoding (OEnc) $h(\cdot)$, relation encoding (REnc) $f_{\{q,k\}}$ to enhance the model’s understanding on neighbor parts, and symmetry encoding (SEnc) $g(\cdot)$ for modeling parts attributes.
Order Encoding. To leverage the assembly pattern in parts, order encoding reveals order information to transformers: $h(v_{i},i)=v_{i}\oplus e_{N}(i)$, where $\oplus$ denotes feature concatenation and $e_{N}(i)$ is a one-hot embedding with the $i$-th position being 1 in a vector of length $N$. This encoding integrates the $i$-th order information into the part feature $v_{i}$ before passing part feature to the transformers.
@@
where $R_{\Theta,i}^{d}=\text{diag}(A_{1},\cdots,A_{d/2})$, $A_{i}=\begin{bmatrix}\text{cos}m\theta_{i}&-\text{sin}m\theta_{i}\\ \text{sin}m\theta_{i}&\text{cos}m\theta_{i}\end{bmatrix}$ is a rotation matrix, and $\text{diag}(\cdot)$ denotes diagonalizing block matrices. The relation encoding is applied to query and key to captures relative order information:
@@
Symmetry Encoding. Symmetrical relations, such as translational (e.g. identical and parallel arms of a chair), reflective (e.g. mirror symmetrical arms of a chair), and rotational symmetry (e.g. identical but not parallel legs of a chair), are proven beneficial for creating structured objects in generative methods [31]. To incorporate symmetrical relations into part features, we use principal component axes via PCA to acqurie 3D bounding boxes and calculate bounding box sizes, which is invariant to input poses. The parts are organized based on identical bounding box sizes, resulting in the partition $\mathcal{P}=\{\mathcal{P}_{1},\cdots,\mathcal{P}_{M}\}$, where each part $p_{i}$ is associated with a unique group $s_{i}\in{1,\cdots,M}$. Symmetry encoding explicitly models symmetrical information into part features: $g(v_{i})=v_{i}\oplus e_{N}(s_{i})$.

### v3

@@ Abstract
We introduce SPAFormer, an innovative model designed to overcome the combinatorial explosion challenge in the 3D Part Assembly (3D-PA) task. This task requires accurate prediction of each part’s poses in sequential steps. As the number of parts increases, the possible assembly combinations increase exponentially, leading to a combinatorial explosion that severely hinders the efficacy of 3D-PA. SPAFormer addresses this problem by leveraging weak constraints from assembly sequences, effectively reducing the solution space’s complexity. Since the sequence of parts conveys construction rules similar to sentences structured through words, our model explores both parallel and autoregressive generation. We further strengthen SPAFormer through knowledge enhancement strategies that utilize the attributes of parts and their sequence information, enabling it to capture the inherent assembly pattern and relationships among sequentially ordered parts. We also construct a more challenging benchmark named PartNet-Assembly covering 21 varied categories to more comprehensively validate the effectiveness of SPAFormer. Extensive experiments demonstrate the superior generalization capabilities of SPAFormer, particularly with multi-tasking and in scenarios requiring long-horizon assembly. Code is available at https://github.com/xuboshen/SPAFormer.
Figure 1: We propose SPAFormer, a novel sequence-conditioned transformer-based method to assemble objects from given parts with 3D point clouds. SPAFormer efficiently leverages the part geometry and sequence information, achieving significantly more plausible assemblies on our constructed benchmark PartNet-Assembly than other baseline methods including ScorePA [3], DGL [15], and RGL [37].
00footnotetext: *Qin Jin is the corresponding author.
@@ 1 Introduction
Figure 2: Illustration of the combinatorial explosion challenge inherent in the assembly process. Specifically: (a) For an object composed of $n$ parts, where we assume each part can occupy one of $m$ discrete positions, the potential combinations of these parts grow at an extraordinary rate, exceeding $O(m^{n})$ in complexity. (b) the number of constituent parts increases when the target object for assembly becomes more complex.
@@
There has been strong anticipation for agents to autonomously assemble parts into an intricate structural object [14, 15, 21, 51, 4]. To achieve this goal, 3D part assembly (3D-PA) is essential [15], which requires the ability of an agent to identify, reorient, and connect geometric parts in the 3D space, ultimately forming a desired object. Such capability empowers an agent to undertake a variety of applications in real-world scenarios, ranging from assembling IKEA furniture [51] to crafting designs in Computer-Aided Design (CAD) environments [4].
@@
Due to its critical importance, the 3D-PA task has attracted increasing attention from researchers. Traditionally, many approaches [15, 3, 59] achieve target object assembly by relying solely on part-level information or basic structural knowledge (e.g., object category label) without additional constraints. We consider these approaches as condition-free approaches. However, the 3D-PA task encounters a substantial obstacle, known as “combinatorial explosion” (depicted in Figure 2), where the number of possible combinations rises beyond exponentially with increasing parts. This phenomenon poses a significant challenge and makes condition-free approaches less effective in achieving accurate part assembly, particularly as the assembly sequence lengthens.
@@
To address the combinatorial explosion challenge, several studies [37, 28, 27, 3, 60, 29] have leveraged visual priors of the object, such as point clouds [28] or RGB images [27, 50], to inform the assembly process. These visual-conditioned methods enable the foresight of an object’s final form and overall appearance to guide the assembly process. However, the acquisition of such detailed visual priors incurs potential costs, reducing the practicality of these approaches in real-world applications. As an alternative, recent initiatives explore a more balanced way by utilizing structural prior knowledge. This includes the adoption of a hierarchical structure graph [44, 35, 13, 19, 49] or employing peg-hole joints on parts [29] for guidance. While this strategy decreases the need for detailed visual priors and is generally less burdensome, the reliance on structural priors still implies an additional expense compared to condition-free approaches. Given the limitations of previous works, we propose that leveraging the assembly sequences represents a more viable solution,which is advantageous due to its weak constraints: assembly sequences are readily obtainable from diverse sources, including instructional manuals [51, 50, 58, 30], the analysis of physical disassembly process [45, 32, 42], or through the prediction of assembly sequences themselves [33, 46]. Similar to how the sequence of words in a sentence determines the meaning, the assembly order of parts conveys the rules for object construction. Consequently, recognizing patterns within these sequences can effectively benefit 3D-PA, offering an efficient method to address combinatorial explosion.
@@
In this work, we propose Sequential 3D Part Assembly with Transformers (SPAFormer), a novel transformer-based model to address the combinatorial explosion challenge in the 3D-PA task. SPAFormer conditioned on the assembly sequence enhances the 3D geometric understanding through reasoning relations among object parts. Specifically, we explore the effectiveness of potential transformer paradigms for 3D-PA, including parallel generator and autoregressive generator. Moreover, we employ knowledge enhancement strategies to enrich the model with part attributes and assembly sequence insights, by incorporating critical features such as part symmetry (e.g., the legs of a chair), assembly order, and relative position of the parts. To comprehensively evaluate our model, we expand the scope of the traditional benchmark [15], enriching it from the limited 3 categories (chair, table, lamp) to a wider variety of 21 categories that range from structured furniture (e.g., chairs, tables) to everyday items (e.g., earphones, clocks). Our findings reveal that SPAFormer not only significantly outperforms existing methods that are condition-free or sequence-conditioned (Figure 1), but also achieves competitive performance compared to approaches employing visual priors.
@@
    Innovative Framework: We propose SPAFormer, a transformer-based model for 3D object assembly with sequential parts. Our model particularly leverages sequential part information and incorporates knowledge enhancement strategies to significantly improve the assembly performance.
@@
    Generalization of SPAFormer: SPAFormer shows superior generalization in object assembly from three crucial perspectives: a) category-specific, enabling it to handle various objects in the same category; b) multi-task, showcasing its versatility across diverse object categories using the same model for the first time; and c) long-horizon, proving its ability in managing complex assembly tasks with numerous parts.
@@
    More Comprehensive Benchmark: To facilitate a thorough evaluation of different models, we introduce an extensive benchmark named PartNet-Assembly, covering up to 21 object categories and providing a broad spectrum of object assembly tasks.
@@ 2 Related Works
Assembly-based 3D Modeling. The assembly-based 3D modeling involves creating structural 3D shapes by integrating information from individual parts. Prior works [1, 20, 17, 11] have explored constructing diverse 3D shapes from a given set of part candidates with precise language semantics. With the development of generative methods, some works [53, 12, 35, 24, 13] focus on leveraging generative models, such as variational autoencoders [55, 25], to generate 3D shapes. Although generative models produce aesthetic 3D shapes, they are not able to deal with assembling true objects in the real world. Many works focus on fracture object assemblies [31, 2, 39, 54, 23, 16], where the model aggregates broken pieces into an object. However, the task of restoring broken objects is not applicable to complex furniture or toy assemblies. Different from previous tasks, the autonomous 3D object assembly [15] is more beneficial for robotic assembly [22, 21, 14, 57, 4] where the parts maintain semantics like IKEA assembly.
@@
Autonomous 3D Object Assembly. In the context of 3D object assembly, the objective is to accurately predict the translation and rotation parameters of individual parts to reconstruct the object’s shape. Many studies [15, 59, 52, 3] try to tackle this task by simply utilizing the part information and category-level priors. Notably, Huang et al. [15] employ graph neural networks to facilitate the understanding of interrelations between parts. Additionally, by expanding data through the inclusion of fractured objects, Zhang et al. [60] demonstrate significant improvements in assembling chairs. Despite these advancements, the problem of combinatorial explosion persists. To mitigate this issue, various prior knowledge on objects has been provided as extra information, including ground-truth object image [27], peg-hole joints on parts [29] and object point clouds [28]. However, these constraints may be costly and less feasible in real-world applications. Instead, advancements in sequence-conditioned part assembly, such as RGL [37], have demonstrated the potential of using weakly constrained assembly sequences for exploring free-form assembly. Building upon these insights, our work introduces a transformer-based framework designed to leverage the advantages offered by assembly sequences in the 3D-PA task.
@@
Positional Encoding in Transformers. Positional encoding facilitates transformers to interpret position information, as evidenced across a diverse range of research domains [56, 6, 34, 26, 47, 7, 10]. Traditionally, the literature consists of two primary encodings: absolute and relative. Absolute positional encoding [48] enables transformers to differentiate words through position-dependent embeddings. Nonetheless, it exhibits limitations in representing the order of words relative to one another [40]. To address this issue, relative positional encoding aims to better characterize the relationships between adjacent words. A notable implementation is the RoPE [41], which rotates the input embedding to effectively extract relative positioning. In the 3D-PA task, Li et al. [27] and Narayan et al. [37] attempt to differentiate identical parts using positional encodings, but the unexpected similar features lead to their suboptimal results. Considering these works, we propose to combine variants of positional encodings to identify assembly patterns among sequential parts.
@@ 3 Methodology
Given a set of $N$ object parts $\mathcal{P}=\{p_{i}\}_{i=1}^{N}$, the part assembly task aims to predict their poses $\{y_{i}\}_{i=1}^{N}\subset\mathrm{SE}(3)$, where $p_{i}$ denotes the 3D point cloud of the $i$-th part and ${\rm SE(3)}$ includes all translation and rotation transformations. The resulting assembled 3D shape $S=\bigcup_{i}^{N}y_{i}(p_{i})$ is formulated by transforming each part-level point cloud $p_{i}$ using its corresponding pose $y_{i}$. During the assembly process, the assembly chain $\mathcal{C}=\{(p_{i},p_{j})\mid p_{i}\preceq p_{j},\ \forall i\leq j\}$ is imposed on the set $\mathcal{P}$, where $p_{i}\preceq p_{j}$ denotes that part $p_{i}$ is assembled before $p_{j}$. The goal of our work is sequential part assembly, which predicts the poses $\{y_{i}\}_{i=1}^{N}$ of $\mathcal{P}$ guided by the predefined assembly chain $\mathcal{C}$.
The overall framework of our proposed SPAFormer is illustrated in Figure 3. In the following sections, we first introduce our knowledge enhancement strategies to effectively incorporate part attributes and relation information in Section 3.1. Then we discuss generator variants for better modeling part relationship in Section 3.2. These strategies include order encoding (OEnc) and relation encoding (REnc) which aim to capture assembly sequence patterns, and symmetry encoding (SEnc) which focuses on enriching part features with symmetry information. In Section 3.3, we introduce our training process.
Figure 3: Illustration of overall end-to-end framework of SPAFormer. (a) The shared 3D backbone extracts the geometry feature of individual parts, followed by (b) knowledge enhancement of part features, which incorporates symmetry, order, and relation information into part features through positional encodings, then generates poses by either (c1) parallel generator, which generates poses of all parts at once, or (c2) autoregressive generator, which decodes poses of parts according to assembly sequences step by step.
@@ 3.1 Knowledge Enhancement for Parts
Due to the data deficiency, which can lead to inferior results when using the data-hungry transformers, we employ several strategies to enhance the model’s understanding of parts. These strategies include order encoding (OEnc) $h(\cdot)$, relation encoding (REnc) $f_{\{q,k\}}$ to enhance the model’s understanding of neighbor parts, and symmetry encoding (SEnc) $g(\cdot)$ for modeling parts attributes.
Order Encoding. To leverage the assembly pattern in part sequence, order encoding reveals the order information to transformers: $h(v_{i},i)=v_{i}\oplus e_{N}(i)$, where $\oplus$ denotes feature concatenation and $e_{N}(i)$ is a one-hot embedding with the $i$-th position being 1 in a vector of length $N$. This encoding integrates the $i$-th order information into the part feature $v_{i}$ before passing the part feature to the transformers.
@@
where $R_{\Theta,i}^{d}=\operatorname{diag}(A_{1},\dots,A_{d/2})$, $A_{i}=\begin{bmatrix}\cos(m\theta_{i})&-\sin(m\theta_{i})\\ \sin(m\theta_{i})&\cos(m\theta_{i})\end{bmatrix}$ is a rotation matrix, and $\operatorname{diag}(\cdot)$ denotes diagonalizing block matrices. The relation encoding is applied to queries and keys to capture the relative order information:
@@
Symmetry Encoding. Symmetrical relations, such as translational (e.g., identical and parallel arms of a chair), reflective (e.g., mirror-symmetrical arms of a chair), and rotational symmetry (e.g., identical but not parallel legs of a chair), are proven beneficial for creating structured objects in generative methods [35]. To incorporate symmetrical relations into part features, we use principal component axes via PCA to acquire 3D bounding boxes and calculate bounding box sizes, which are invariant to input poses. The parts are organized based on identical bounding box sizes, resulting in the partition $\mathcal{P}=\{\mathcal{P}_{1},\cdots,\mathcal{P}_{M}\}$, where each part $p_{i}$ is associated with a unique group $s_{i}\in\{1,\dots,M\}$. Symmetry encoding explicitly models symmetrical information into part features: $g(v_{i})=v_{i}\oplus e_{N}(s_{i})$.